1 Hydrogen gas reduces NO to N2 in the following reaction:
2H2 (g) +2 NO (g) ⇌ 2H2O (g) + N2(g)
The initial reaction rates of four mixtures of H2 and NO were measured at 1500°C with the following results:
Experiment [H2]0 (M) [NO]0 (M) Initial Rate (M/s)
1 0.212 0.136 0.0221
2 0.212 0.272 0.0445
3 0.424 0.544 0.349
4 0.848 0.544 1.40
a) what is the rate la for the reaction at 1500°C?
b) Determine the rate constant for the reaction at 1500°C?
1
Expert's answer
2020-03-25T12:48:19-0400

"rate = k [NO]^2 [H2]"

From experiment 1 and 2

"rate=dC\/dT" =0.136/1500= 9.07*10-5

9.07*10-5= K(0.272)2(0.212)

K=5.78M-2S-1